OpenStudy (anonymous):

0=5/6x+2/7x+1 when i trien to solve this two different ways once i got x=-1 and the other way i got x=-1/47 both are wrong help please

OpenStudy (barrycarter):

Well, how did you try to solve it?

OpenStudy (anonymous):

Clarification on the equation. \[0=\frac{5}{6}x+\frac{2}{7}x+1? \space or \space 0=\frac{5}{6x}+\frac{2}{7x}+1?\]

OpenStudy (anonymous):

cliffsedge its the first one you wrote. how did you get that typed.

OpenStudy (anonymous):

The reply box has options for Equation, Draw, and Attach File.

OpenStudy (anonymous):

barrycarter i tried multiplying both sides by 42 to get rid of the fractions

OpenStudy (barrycarter):

OK, what happens when you multiply by 42 (good technique)

OpenStudy (anonymous):

ny reply box has none of those. i wonder if it is because i am using an ipad

OpenStudy (anonymous):

after multiplying by 42 0=35x+12x+42

OpenStudy (barrycarter):

Looks good so far... what next?

OpenStudy (anonymous):

0=42x+42

OpenStudy (barrycarter):

Wait, what's 35+12 ?

OpenStudy (anonymous):

haaaaa holy cow. i am an idiot sorry

OpenStudy (anonymous):

Moral of the story: The hardest part about algebra is the arithmetic.

OpenStudy (anonymous):

yes thank you i am embarrassed

OpenStudy (anonymous):

Happens to everybody - just a little earlier I factored x^2-4 as (x+4)(x-4). :">